OpenStudy (anonymous):

If the sin 60° = square root of three over two, then which statement is true?

OpenStudy (anonymous):

cos 30° = square root of three over two, because the cosine and sine are complements cos 120° = square root of three over two, because the cosine and sine are supplements cos 30° = 0, because the cosine and sine are complements cos 120° = 0, because the cosine and sine are supplements

OpenStudy (baru):

sin and cos are complemntary which means \(cos(\theta) = sin(90-\theta) \)

OpenStudy (baru):

so which do u think the answer is?

OpenStudy (anonymous):

don't quite understand

OpenStudy (baru):

just substitute \( \theta=60^0 \) and see what u get

OpenStudy (baru):

* substitute in the formula i gave

OpenStudy (anonymous):

so cos 30=0

OpenStudy (anonymous):

is that right or is it cos 30= sr3/2

OpenStudy (baru):

did u try substituting in the formula like i said?

OpenStudy (anonymous):

yes i did i think the answer is a, cos 30= sr3/2 i that right?

OpenStudy (baru):

yes

OpenStudy (baru):

i hope u understood why thats the right answer :)

OpenStudy (anonymous):

yes i do thank you